For the following reaction, Keq is 0.00183 at 390. K:

PCl5(g)  PCl3(g) + Cl2(g)

If 3.72 g of PCl5 is placed in a 3.68 L bulb at 390. K, what is the equilibrium pressure of Cl2? 1 atm = 1.013 bar

bar

The number of significant digits is set to 3; the tolerance is +/-4%
Answer #1

\bf \text{The no. moles of } PCl_5=\frac{3.72\;g}{208.24\;g/mol.}=0.018 \;mol

\bf \text{ Volume of container }=3.68 \;L

\bf PCl_5(g)\rightleftharpoons PCl_3(g) + Cl_2(g)\\ I\;\;\;\;\;0.018\;\;\;\;\;\;\;\;\;\;\;\;0\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;0\\ C\;\;\;\;-x\;\;\;\;\;\;\;\;\;\;\;\;\;\;x\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;x\\ E\;\;0.018-x\;\;\;\;\;\;\;x\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;x\\

Total no. of moles at equilibrium 0.018 + x

\bf \text{ The equilibrium pressure of chlorine }P_{Cl_2}=\frac{x}{0.018+x}\times(1\;atm)

\bf K_{eq}=\frac{[PCl_3][Cl_2]}{[PCl_5]}

\bf 0.00183=\frac{(x/3.68L)(x/3.68L)}{(0.018-x)/3.68\;L}

\bf 0.00183=\frac{x^2}{(0.018-x)\times3.68}

\bf x^2=0.00183\times(0.018-x)\times3.68

\bf x^2=0.0067344\times(0.018-x)

\bf x^2+0.0067344x+0.000121=0

\bf x=0.00814

Total no. of moles at equilibrium 0.018 + x

\bf \text{ Total no. of moles at equilibrium }=0.018+0.00814=0.02614

\bf \text{ The equilibrium pressure of chlorine }P_{Cl_2}=\frac{x}{0.018+x}\times(1\;atm)

\bf \text{ The equilibrium pressure of chlorine }P_{Cl_2}=\frac{0.00814}{0.02614}\times(1\;atm)

\large \bf \text{ The equilibrium pressure of chlorine }P_{Cl_2}=0.312\;atm